OpenStudy (anonymous):

Help understand radival equations? When i solve x+6=radical(x+12) on the calculator.. i get -3 which is a solution that works. Howevet if i square both sides and solve it algebraically. Why do i get a extraneous solution of -8?

OpenStudy (solomonzelman):

you mean, \(\large\color{black}{ x+6=\sqrt{x+12} }\) ?

OpenStudy (solomonzelman):

yes, -3 is the correct solution.

OpenStudy (solomonzelman):

but what is your question then?

OpenStudy (solomonzelman):

oh, I see, you want to actually work the solutions, right?

OpenStudy (anonymous):

Im just wondee why i get anothet solution when i do it algebraically

OpenStudy (anonymous):

Cause i get -8 and when i plug it in to the original problem i get 2=-2 which means -8 is not a solution

OpenStudy (loser66):

The right hand side must be \(\geq 0\), that is the left hand side must be \(\geq 0\) also. To have it, solve \(x+6\geq 0\rightarrow x\geq -6\), That allows you reject x=-8 , since \(-8 < -6\)

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

Wait, but why does squaring it give another solution?

OpenStudy (loser66):

Nope, it gives you 2 solutions which are -3 and -8. You checked and knew -8 is not a solution but you don't know how to get rid of. My logic gives you a key.